rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#3853 — FS#7795 —Optical loop Paris little ring

Attached to Project— ADSL / SDSL
Maintenance
Backend / Core
CLOSED
100%
We have a problem on a fiber. The securisation of some NRA
is no longer active.

For us the problem is between CPY94 and GRA94.

An intervention is being processed.
Date:  Friday, 21 December 2012, 01:40AM
Reason for closing:  Done
Additional comments about closing:  The team has finished the melting. Everything is secured from our side.
Comment by OVH - Friday, 21 December 2012, 00:02AM

The constraint on the fiber has not been found yet.
Many difficulties on the field (Clogged or tarred views)
are making the progression so slow.

In new updates, the constraint will be on a portion
of fiber and not on a new connection.


Comment by OVH - Friday, 21 December 2012, 00:05AM

The constraint has been found fixed for OVH, the buckle is again secured. We have lost 1.1 dB thereafter the melting, but this does not impact the good performance of the buckle.

The team on site are melting the fibers of other customers. We are waiting the confirmation of the operator before the closure of the incident.